# Analytical results for the distribution of first return times of non-backtracking random walks on configuration model networks

### v1

@@ 3 Application to specific random networks
In this section we apply the general results derived above to NBWs on specific random networks that belong to the class of configuration model networks [33]. The configuration model network ensemble is a maximum entropy ensemble under the condition that the degree distribution $P(k)$ is imposed. To generate a network instance drawn from an ensemble of configuration model networks consisting of $N$ nodes with a given degree distribution $P(k)$, one draws the degrees of the $N$ nodes independently from $P(k)$. This gives rise to a degree sequence of the form $k_{1},k_{2},\dots,k_{N}$. Each node $i$ is then attached to $k_{i}$ half edges or stubs. The construction proceeds by connecting, at each step, a random pair of stubs that belong to two different nodes $i$ and $j$ that are not already connected, forming an edge between them. This procedure is repeated until all the stubs are exhausted. The process may get stuck before completion in case that all the remaining stubs belong to the same node or to pairs of nodes that are already connected. In such case one needs to perform some random reconnections in order to complete the construction.
In particular we study the first return process on RRGs, ER networks and configuration model networks with exponential and power-law degree distributions. For each type of network we calculate the tail distribution of first return times as well as the mean and variance.
@@ 3.1 Random regular graphs
Random regular graphs are random networks of a finite size in which all the nodes are of the same degree, but the connectivity is random [8]. They thus belong to the class of configuration model networks. Consider an RRG that consists of $N$ nodes of degree $c\geq 3$. In such network, all the nodes reside on a single connected component. As a result, an RW (or NBW) starting from any initial node $i$ may reach any other node $j$.
@@
It would be useful to compare the distribution of first return times of NBWs on RRGs to the corresponding distribution of classical RWs on RRGs. The latter distribution consists of a contribution from retroceding trajectories, which are dominant at short times and non-retroceding trajectories, which are dominant at long times. The distribution $P(T_{\rm FR}>t)$, given by Eq. (29), is analogous to the contribution of the non-retroceding RW trajectories in classical RWs, which for sufficiently long times is given by [7]
@@
In Fig. 1 we present analytical results for the tail distribution $P(T_{\rm FR}>t)$ (solid lines) of first return times of an NBW on an RRG of size $N=1000$ and degree $c=10$. The analytical results, obtained from Eq. (29), are in very good agreement with the results obtained from computer simulations (circles).
Figure 1: Analytical results for the tail distribution $P(T_{\rm FR}>t)$ (solid line) of first return times of an NBW on an RRG of size $N=1000$ and degree $c=10$. The analytical results, obtained from Eq. (29), are in very good agreement with the results obtained from computer simulations (circles).
@@
thus the mean first return time does not depend on the degree $c$. This result is in agreement with Kac’s lemma, expressed by Eq. (1). Since all the nodes in an RRG are of the same degree, the probability that an RW (or an NBW) will reside at any given node at time $t$ is $P_{i}(\infty)=1/N$. Inserting $P_{i}(\infty)$ into Eq. (1), we obtain ${\mathbb{E}}[T_{\rm FR}]=N$. This result is also in agreement with the mean first return time of an RW on an RRG, calculated in Ref. [7].
@@
Interestingly, for a classical RW on an RRG, it was found that the variance of the distribution of first return times is given by [7]
@@
This result is larger than the variance for NBWs by a multiplicative factor of $\frac{c}{c-2}$. This factor is significant for sparse RRGs and approaches $1$ as $c$ is increased. It is due to the fact that in classical RWs the distribution of first return times is bimodal, consisting of two different types of first return trajectories. At short times it is dominated by retroceding trajectories while at long times it is dominated by non-retroceding trajectories. This separation of time scales broadens the distribution and increases the variance. The difference in the variance between NBWs and classical RWs reflects the fact that the Kac’s lemma applies only to the mean first return time and does not provide any prediction for the variance.
@@ 3.2 Erdős-Rényi networks
for $k=0,1,2,\dots$, where $c=(N-1)p$ is the mean degree $\langle K\rangle$ and the variance is given by ${\rm Var}(K)=c$.
In general, for $c>1$ an ER network consists of a giant component and finite tree components. Since we focus in this paper on networks that consist of a single connected component, we restrict ourselves to the case in which $c>\ln N$, where in the large network limit the giant component encompasses the whole network [8]. In the case that $c>\ln N$, the probability that a random node will be isolated is $P(K=0)<1/N$, which implies that in a typical network instance the expected number of isolated nodes will be smaller than $1$. Since we study NBWs we would like to ensure that the network instances we consider will also not include leaf nodes of degree $k=1$. Therefore, in the analysis we focus on the limit of sufficiently dense networks that satisfy $c>-W(-1/N)$, where $W(x)$ is the Lambert W function [39]. In this limit the probability that a random node will be a leaf node satisfies $P(K=1)<1/N$. In practice, when we generate network instances for the computer simulations, we discard network instances that include isolated nodes or leaf nodes.
@@
Figure 2: Analytical results for the tail distribution $P(T_{\rm FR}>t|K>1)$ (solid line) of first return times of an NBW on an Erdős-Rényi network of size $N=1000$ and mean degree $c=10$. The analytical results, obtained from Eq. (38), are in very good agreement with the results obtained from computer simulations (circles).
@@
Figure 3: Analytical results for the mean first return time ${\mathbb{E}}[T_{\rm FR}|K>1]$ (solid line) of an NBW on an Erdős-Rényi network of size $N=1000$, as a function of the mean degree $c$, for $c>\ln N$, where the whole network consists of a single connected component and network instances that include leaf nodes are discarded. The analytical results, obtained from Eq. (41), are in very good agreement with the results obtained from computer simulations (circles).
@@
Carrying out the summation on the right hand size of Eq. (44), we obtain
@@
where $\,{}_{3}F_{3}\Bigl{(}\Big{.}\begin{array}[]{c}a_{1},a_{2},a_{3}\\ b_{1},b_{2},b_{3}\end{array}\Big{|}z\Bigr{)}$ is the generalized hypergeometric function [39]. Thus, the variance of $P(T_{\rm FR}>t|K>1)$ is give by
@@
Figure 4: Analytical results for the variance ${\rm Var}(T_{\rm FR}|K>1)$ (solid line) of the distribution of first return times of an NBW on an Erdős-Rényi network of size $N=1000$, as a function of the mean degree $c$. The analytical results, obtained from Eq. (49), are in very good agreement with the results obtained from computer simulations (circles).
@@ 3.3 Configuration model networks with an exponential degree distribution
where $\alpha>0$ is the rate parameter and the degree $k$ takes values in the range $k_{\rm min}\leq k\leq\infty$ [$P(k)=0$ for $10\leq k\leq k_{\rm min}-$ 1]. The parameter $A$ is a normalization factor and it is given by $A=(1-e^{-\alpha})e^{\alpha k_{\rm min}}$. In order to obtain a network that consists of a single connected component, one needs to choose $k_{\rm min}\geq 2$.
Imposing the normalization condition and parametrizing the distribution in terms of the mean degree $c=\langle K\rangle$, one can rewrite the degree distribution in the form
@@
Figure 5: Analytical results for the tail distribution $P(T_{\rm FR}>t)$ (solid lines) of first return times of an NBW on a configuration model network of size $N=1000$ which exhibits an exponential degree distribution with $k_{\rm min}=3$ and mean degree $c=10$. The analytical results, obtained from Eq. (56), are in very good agreement with the results obtained from computer simulations (circles).
@@
where $H_{m}$ is the Harmonic number [39]. Note that the leading term in ${\mathbb{E}}[T_{\rm FR}]$ is given by $N\ln c$, unlike the the RRG and ER network, in which ${\mathbb{E}}[T_{\rm FR}]\simeq N$. This reflects the fact that the variance of the degrees in the exponential case is much larger than in the Poisson distribution and that it increases as $c$ is increased. Since the exponential distribution is broad and highly asymmetric, the expansion presented in Eq. (22) cannot be used to reproduce the results of Eq. (59). Eq. (59) shows that the mean first return time in configuration model networks with an exponential degree distribution increases logarithmically with the mean degree $c$.
@@
Figure 6: Analytical results for the mean first return time ${\mathbb{E}}[T_{\rm FR}]$ (solid line) of an NBW on a configuration model network of size $N=1000$, which exhibits an exponential degree distribution with $k_{\rm min}=3$, as a function of the mean degree $c$. The analytical results, obtained from Eq. (58), are in very good agreement with the results obtained from computer simulations (circles).
@@
Figure 7: Analytical results for the variance ${\rm Var}(T_{\rm FR})$ (solid line) of the distribution of first return times of an NBW on a configuration model network of size $N=1000$, which exhibits an exponential degree distribution with $k_{\rm min}=3$, as a function of the mean degree $c$. The analytical results, obtained from Eq. (62), are in very good agreement with the results obtained from computer simulations (circles).
@@ 3.4 Configuration model networks with a power-law degree distribution
Since a power-law distribution may allow nodes of high degree, it is important to note that in order to enable the construction of a configuration model network with no degree-degree correlations the maximal degree should satisfy $k_{\rm max}<\sqrt{N}$. This condition should be maintained when one takes the limit of large $k_{\rm max}$, which implies that the network size $N$ must be increased accordingly.
The mean degree is given by
@@
For $\gamma\leq 2$ the mean degree (and the variance) diverge when $k_{\rm max}\rightarrow\infty$. For $\gamma>3$ both the mean degree and the variance are bounded. In the intermediate range of $32<\gamma<$ 3 the mean degree $\langle K\rangle$ is bounded while the variance ${\rm Var}(K)$ diverges. In this regime, as $k_{\rm max}$ is increased, the variance diverges like
@@
Figure 8: Analytical results for the tail distribution $P(T_{\rm FR}>t)$ (solid lines) of first return times of an NBW on a configuration model network of size $N=1000$ which exhibits a power-law distribution with $k_{\rm min}=3$, $k_{\rm max}=30$ and $\gamma=2.5$. The analytical results, obtained from Eq. (68), are in very good agreement with the results obtained from computer simulations (circles).
@@
[…]
where $\Gamma(y)$ is the Gamma function [39]. In the context of this paper the Hurwitz zeta function $\zeta(\gamma,k)$ is evaluated in the range of $\gamma>1$ and $k\geq 3$. Exploring the terms on the right hand size of Eq. (72) in this range of values, it was found that the contribution of the integral is negligible. Thus, Eq. (72) can be simplified to
@@ 4 Discussion
A key observation, expressed by Eq. (10) is that the distribution of first return times for an initial node of a given degree $k$ depends only on the degree $k$ and on the total number of ’directed’ edges in the network, given by $Nc$. It does not depend on the degree distribution $P(k)$, which accounts for the way in which the $Nc-2k$ remaining ’directed’ edges are divided among the $N-1$ remaining nodes. This implies that the distribution of first return times is determined by local properties of the network and is not sensitive to the global structure.
@@
In Fig. 9 we present analytical results for the distribution of first return times of NBWs starting from a random initial node of degree $k=8$ in a configuration model network of size $N=1000$ and mean degree $c=8$, given by Eq. (10). The analytical results are found to be in very good agreement with simulation results for RRGs ($\circ$), ER networks ($\times$) and configuration model networks with exponential ($+$) and power-law ($\square$) distributions.
@@
Figure 9: Analytical results for the distribution of first return times for NBWs starting from an initial node of degree $k=8$ in configuration model networks of size $N=1000$ and mean degree $c=8$, given by Eq. (10). The analytical results are found to be in very good agreement with simulation results for RRGs ($\circ$), ER networks ($\times$) and configuration model networks with exponential ($+$) and power-law ($\square$) distributions.
@@
Comparing the results obtained for the four random network models considered above, we conclude that the mean first return time strongly depends on the variability of the degrees of nodes in the network. More specifically, as the degree distribution $P(k)$ becomes broader the mean first return time ${\mathbb{E}}[T_{\rm FR}]$ increases. This is illustrated by the fact that for an NBW on an RRG ${\mathbb{E}}[T_{\rm FR}]\simeq N$, for an NBW on an ER network ${\mathbb{E}}[T_{\rm FR}|K>1]\simeq N\left(1+\frac{1}{c}\right)$, for an NBW on a configuration model network with an exponential degree distribution ${\mathbb{E}}[T_{\rm FR}]\simeq N\ln c$ and for an NBW on a configuration model network with a power-law degree distribution ${\mathbb{E}}[T_{\rm FR}]\sim Nc/k_{\rm min}$. In light of these results, it is interesting to note that the dependence of the mean first return time on the mean degree $c$ is a non-trivial issue, which depends on the details of the degree distribution. In the examples studied here we observe three different behaviors: in RRGs the mean first return time is indepndent of $c$, in ER networks it decreases with $c$ and in configuration model networks with an exponential degree distribution it increases with $c$.
@@
In all the network ensembles considered above, the long-time tail of $P(T_{\rm FR}>t)$ exhibits a decaying exponential form, which is determined by the lowest-degree nodes in the network. From a broader perspective, it implies that the distribution of first return times is mostly characterized by low-degree nodes that reside in the periphery of the network. This is unlike the outburst dynamics of other processes such as the spreading of information and infections, which are dominated by the highest degree nodes (or hubs) that reside in the core of the network [5].
@@
Apart from the first return process, there are other significant events that take place over the lifetime of an NBW (and other RWs) on a random network. One of them is the first hitting (FH) process, which is the first time at which an NBW steps into a previously visited node. Starting from a random initial node $i$, in the early stages of its trajectory, an NBW visits a new node at each time step. During this time, the statistical properties of the NBW trajectory are identical to those of a self avoiding walk [45]. After the first hitting event, in some of the time step the NBW visits yet-unvisited nodes and in other time steps it revisits nodes that it has already visited before. The distribution of first hitting times of RWs and NBWs on ER networks were studied in Refs. [43] and [44], respectively. It was found that in both cases, for sufficiently dense ER networks in which there are no leaf nodes of degree $k=1$, the distribution $P(T_{\rm FH}>t)$ of first hitting times is given by a product of an exponential distribution and a Rayleigh distribution, which is a special case of the Weibull distribution. In this limit, the mean first hitting time of NBWs on ER networks is given by
@@
The results presented in this paper were derived in the context of configuration model networks. However, we expect them to apply within a good approximation to a somewhat broader class of small-world networks which are sufficiently strongly connected without bottlenecks and exhibit short mixing times of the NBW, determined by the spectral gap of the Hashimoto matrix. In contrast, these results are not expected to apply in the case of modular networks, which consist of several models with weak connections between them and for networks that exhibit long mixing times. However, in the case of modular networks, at short times, the distribution of first return times is likely to behave as if the module on which the initial node resides is isolated from other modules. This behavior persists until the probability that the NBW will hop into some other module becomes significant.
@@
Another key factor influencing the mixing times is the clustering coefficient, primarily mediated through its effect on the spectral gap. The clustering coefficient measures how often a node’s neighbors form triangles indicating the degree of local inter-connectedness. Networks with higher clustering coefficients typically have smaller spectral gaps. This occurs because increased clustering introduces more local structure. A smaller spectral gap leads to longer mixing times, since random walks before trapped in tightly connected netighborhood before fully exploring the network [53]. Consequently, higher clustering increases the mixing times. We thus expect our results to be valid as long as the clustering is not too strong.

### v2

@@ 5 Application to specific random networks
In this section we apply the general results derived above to NBWs on specific random networks that belong to the class of configuration model networks [32]. In particular we study the first return process on RRGs, ER networks and configuration model networks with exponential and power-law degree distributions. For each type of network we calculate the tail distribution of first return times as well as the mean and variance.
@@ 5.1 Random regular graphs
Random regular graphs are random networks of a finite size in which all the nodes are of the same degree, but the connectivity is random [10]. They thus belong to the class of configuration model networks. Consider an RRG that consists of $N$ nodes of degree $c\geq 3$. In such network, in the large $N$ limit, all the nodes reside on a single connected component. As a result, an RW (or NBW) starting from any initial node $i$ may reach any other node $j$.
@@
It would be useful to compare the distribution of first return times of NBWs on RRGs to the corresponding distribution of simple RWs on RRGs. The latter distribution consists of a contribution from retroceding trajectories, which are dominant at short times and non-retroceding trajectories, which are dominant at long times. The distribution $P(T_{\rm FR}>t)$, given by Eq. (32), is analogous to the contribution of the non-retroceding RW trajectories in simple RWs, which for sufficiently long times is given by [9]
@@
In Fig. 1 we present analytical results, obtained from Eq. (32), for the tail distribution $P(T_{\rm FR}>t)$ (solid line) of first return times of an NBW on an RRG of size $N=1000$. Note that the right hand side of Eq. (32) does not depend on the degree $c$, which implies that the results are valid for RRGs with any degree $c\geq 3$. Indeed, the analytical results are found to be in very good agreement with the results obtained from computer simulations for RRGs with $c=3$ ($\times$) and $c=10$ ($\circ$).
Figure 1: Analytical results (solid line), obtained from Eq. (32), for the tail distribution $P(T_{\rm FR}>t)$ of first return times of an NBW on an RRG of size $N=1000$. The right hand side of Eq. (32) does not depend on the degree $c$, which implies that these results are valid for RRGs with any degree $c\geq 3$. Indeed, the analytical results are in very good agreement with the results obtained from computer simulations for $c=3$ ($\times$) and for $c=10$ ($\circ$). Each data point of the simulation results was obtained by averaging the results obtained for 2020 network instances and $100,000$ NBW trajectories for each network instance.
The mixing time of an NBW on an RRG of size $N$ and degree $c$ is given by [54]
$t_{\rm mix}(N,c)=\frac{\ln N}{\ln(c-1)}+{\mathcal{O}}(1).$ (34)
Applying this result to NBWs on RRGs of size $N=1000$ with degrees of $c=3$ and $c=10$, it is found that $t_{\rm mix}(1000,3)\simeq 10$ and $t_{\rm mix}(1000,10)\simeq 3$, which are clearly much smaller than the time scales that are relevant to the first return process.
For the simulations we generated $20$ independent instances of the network. On each network instance, we generated $100,000$ NBW trajectories, where each trajectory starts from a random initial node $i$ at time $t=0$. Each NBW trajectory was terminated upon its first return to the initial node $i$. The first return time $t$ is thus equal to the length of the trajectory. The simulation results were obtained by averaging the results over all these trajectories.
@@
thus the mean first return time does not depend on the degree $c$. This result is in agreement with Kac’s lemma, expressed by Eq. (1). Since all the nodes in an RRG are of the same degree, the probability that an RW (or an NBW) will reside at any given node at time $t$ is $P_{i}(\infty)=1/N$. Inserting $P_{i}(\infty)$ into Eq. (1), we obtain ${\mathbb{E}}[T_{\rm FR}]=N$. This result is also in agreement with the mean first return time of a simple RW on an RRG, calculated in Ref. [9].
@@
Interestingly, for a simple RW on an RRG, it was found that the variance of the distribution of first return times is given by [9]
@@
This result is larger than the variance for NBWs by a multiplicative factor of $\frac{c}{c-2}$. This factor is significant for sparse RRGs and approaches $1$ as $c$ is increased. It is due to the fact that in simple RWs the distribution of first return times is bimodal, consisting of two different types of first return trajectories. At short times it is dominated by retroceding trajectories while at long times it is dominated by non-retroceding trajectories. This separation of time scales broadens the distribution and increases the variance. The difference in the variance between NBWs and simple RWs reflects the fact that Kac’s lemma applies only to the mean first return time and does not provide any prediction for the variance.
@@ 5.2 Erdős-Rényi networks
for $k=0,1,2,\dots$, where c=(N1)pc=(N-1)p is the mean degree $\langle K\rangle$ and the variance is given by ${\rm Var}(K)=c$.
In general, for $c>1$ an ER network consists of a giant component and finite tree components. Since we focus in this paper on networks that consist of a single connected component, we restrict ourselves to the case in which $c>\ln N$, where in the large network limit the giant component encompasses the whole network [10, 6]. In the case that $c>\ln N$, the probability that a random node will be isolated is $P(K=0)<1/N$, which implies that in a typical network instance the expected number of isolated nodes will be smaller than $1$. Since we study NBWs we would like to ensure that the network instances we consider will also not include leaf nodes of degree $k=1$. Therefore, in the analysis we focus on the limit of sufficiently dense networks that satisfy c>W(1/N)c>-W(-1/N), where $W(x)$ is the Lambert W function [70]. In this limit the probability that a random node will be a leaf node satisfies $P(K=1)<1/N$. In practice, when we generate network instances for the computer simulations, we discard network instances that include isolated nodes or leaf nodes.
@@
Figure 2: Analytical results for the tail distribution $P(T_{\rm FR}>t|K>1)$ (solid line) of first return times of an NBW on an Erdős-Rényi network of size $N=1000$ and mean degree $c=10$. The analytical results, obtained from Eq. (42), are in very good agreement with the results obtained from computer simulations (circles). The simulation results were obtained using the same averaging procedure as in Fig. 1.
@@
Figure 3: Analytical results for the mean first return time ${\mathbb{E}}[T_{\rm FR}|K>1]$ (solid line) of an NBW on an Erdős-Rényi network of size $N=1000$, as a function of the mean degree $c$, for $c>\ln N$, where the whole network consists of a single connected component and network instances that include leaf nodes are discarded. The analytical results, obtained from Eq. (45), are in very good agreement with the results obtained from computer simulations (circles). Each data point of the simulation results was obtained by averaging the results obtained for 2020 network instances and $10,000$ NBW trajectories for each network instance.
@@
Carrying out the summation on the right hand side of Eq. (48), we obtain
@@
where $_{3}F_{3}\Bigl(\Big.\begin{array}[]{c}a_{1},a_{2},a_{3}\\ b_{1},b_{2},b_{3}\end{array}\Big|z\Bigr)$ is the generalized hypergeometric function [70]. Thus, the variance of $P(T_{\rm FR}>t|K>1)$ is given by
@@
Figure 4: Analytical results for the variance ${\rm Var}(T_{\rm FR}|K>1)$ (solid line) of the distribution of first return times of an NBW on an Erdős-Rényi network of size $N=1000$, as a function of the mean degree $c$. The analytical results, obtained from Eq. (53), are in very good agreement with the results obtained from computer simulations (circles). The simulation results were obtained using the same averaging procedure as in Fig. 3.
@@ 5.3 Configuration model networks with an exponential degree distribution
where $\alpha>0$ is the rate parameter and the degree $k$ takes values in the range $k_{\rm min}\leq k\leq\infty$ [$P(k)=0$ for $0\leq k\leq k_{\rm min}-1$]. The parameter $A$ is a normalization factor and it is given by $A=(1-e^{-\alpha})e^{\alpha k_{\rm min}}$. In order to obtain a network that consists of a single connected component, one needs to choose $k_{\rm min}\geq 2$.
Imposing the normalization condition and parameterizing the distribution in terms of the mean degree $c=\langle K\rangle$, one can rewrite the degree distribution in the form [42]
@@
To explore the asymptotic long time tail of $P(T_{\rm FR}>t)$ we expand the right hand side of Eq. (60) in powers of $\exp\left(-\frac{t}{Nc}\right)\ll 1$. We obtain
@@
Figure 5: Analytical results for the tail distribution $P(T_{\rm FR}>t)$ (solid lines) of first return times of an NBW on a configuration model network of size $N=1000$ which exhibits an exponential degree distribution with $k_{\rm min}=3$ and mean degree $c=10$. The analytical results, obtained from Eq. (60), are in very good agreement with the results obtained from computer simulations (circles). The simulation results were obtained using the same averaging procedure as in Fig. 1.
@@
where $H_{m}$ is the Harmonic number [70]. Note that the leading term in ${\mathbb{E}}[T_{\rm FR}]$ is given by $N\ln c$, unlike the RRG and ER network, in which ${\mathbb{E}}[T_{\rm FR}]\simeq N$. This reflects the fact that the variance of the degrees in the exponential case is much larger than in the Poisson distribution and that it increases as $c$ is increased. Since the exponential distribution is broad and highly asymmetric, the expansion presented in Eq. (25) cannot be used to reproduce the results of Eq. (63). Eq. (63) shows that the mean first return time in configuration model networks with an exponential degree distribution increases logarithmically with the mean degree $c$.
@@
Figure 6: Analytical results for the mean first return time ${\mathbb{E}}[T_{\rm FR}]$ (solid line) of an NBW on a configuration model network of size $N=1000$, which exhibits an exponential degree distribution with $k_{\rm min}=3$, as a function of the mean degree $c$. The analytical results, obtained from Eq. (62), are in very good agreement with the results obtained from computer simulations (circles). The simulation results were obtained using the same averaging procedure as in Fig. 3.
@@
Figure 7: Analytical results for the variance ${\rm Var}(T_{\rm FR})$ (solid line) of the distribution of first return times of an NBW on a configuration model network of size $N=1000$, which exhibits an exponential degree distribution with $k_{\rm min}=3$, as a function of the mean degree $c$. The analytical results, obtained from Eq. (66), are in very good agreement with the results obtained from computer simulations (circles). The simulation results were obtained using the same averaging procedure as in Fig. 3.
@@ 5.4 Configuration model networks with a power-law degree distribution
Since a power-law distribution may allow nodes of high degree, it is important to note that in order to enable the construction of a configuration model network in which degree-degree correlations are negligible, one must impose an upper cutoff on the degree distribution, which satisfies $k_{\rm max}<\sqrt{Nc}$ [38, 39].
The mean degree is given by [42]
@@
For $\gamma\leq 2$ the mean degree (and the variance) diverge when $k_{\rm max}\rightarrow\infty$. For $\gamma>3$ both the mean degree and the variance are bounded. In the intermediate range of $2<\gamma<3$ the mean degree $\langle K\rangle$ is bounded while the variance ${\rm Var}(K)$ diverges. In this regime, as $k_{\rm max}$ is increased, the variance diverges like
@@
Figure 8: Analytical results for the tail distribution $P(T_{\rm FR}>t)$ (solid lines) of first return times of an NBW on a configuration model network of size $N=1000$ which exhibits a power-law distribution with $k_{\rm min}=3$, $k_{\rm max}=30$ and $\gamma=2.5$. The analytical results, obtained from Eq. (72), are in very good agreement with the results obtained from computer simulations (circles). The simulation results were obtained using the same averaging procedure as in Fig. 1.
@@
$\zeta(\gamma,k)=\frac{k^{-\gamma}}{2}+\frac{k^{1-\gamma}}{\gamma-1}+\frac{1}{\Gamma(\gamma)}\int_{0}^{\infty}\left(\frac{1}{e^{x}-1}-\frac{1}{x}+\frac{1}{2}\right)x^{\gamma-1}e^{-kx}dx,$ (76)
where $\Gamma(y)$ is the Gamma function [70]. In the context of this paper the Hurwitz zeta function $\zeta(\gamma,k)$ is evaluated in the range of $\gamma>1$ and $k\geq 3$. Exploring the terms on the right hand side of Eq. (76) in this range of values, it was found that the contribution of the integral is negligible. Thus, Eq. (76) can be simplified to
@@ 6 Discussion
A key observation, expressed by Eq. (13) is that the distribution of first return times for an initial node of a given degree $k$ depends only on the degree $k$ and on the total number of ’directed’ edges in the network, given by $Nc$. It does not depend on the degree distribution $P(k)$, which accounts for the way in which the Nc2kNc-2k remaining ’directed’ edges are divided among the N1N-1 remaining nodes. This implies that the distribution of first return times is determined by local properties of the network and is not sensitive to the global structure.
@@
In Fig. 9 we present analytical results for the distribution of first return times of NBWs starting from a random initial node of degree $k=8$ in a configuration model network of size $N=1000$ and mean degree $c=8$, given by Eq. (13). The analytical results are found to be in very good agreement with simulation results for RRGs ($\circ$), ER networks ($\times$) and configuration model networks with exponential (++) and power-law ($\square$) distributions. As can be seen, the distribution $P(T_{\rm FR}>t|K=8)$ does not depend on the degree distribution $P(k)$ but only on the mean degree $c$ and on the degree of the initial node.
@@
Figure 9: Analytical results for the distribution of first return times for NBWs starting from an initial node of degree $k=8$ in configuration model networks of size $N=1000$ and mean degree $c=8$, given by Eq. (13). The analytical results are found to be in very good agreement with simulation results for RRGs ($\circ$), ER networks ($\times$) and configuration model networks with exponential (++) and power-law ($\square$) distributions. Each data point in the simulation results was obtained by averaging over 2020 independent network instances and $10,000$ NBW trajectories in each network instance, starting from the same initial node of degree $k=8$.
@@
Comparing the results obtained for the four random network models considered above, we conclude that the mean first return time strongly depends on the variability of the degrees of nodes in the network. More specifically, as the degree distribution $P(k)$ becomes broader the mean first return time ${\mathbb{E}}[T_{\rm FR}]$ increases. This is illustrated by the fact that for an NBW on an RRG ${\mathbb{E}}[T_{\rm FR}]\simeq N$, for an NBW on an ER network ${\mathbb{E}}[T_{\rm FR}|K>1]\simeq N\left(1+\frac{1}{c}\right)$, for an NBW on a configuration model network with an exponential degree distribution ${\mathbb{E}}[T_{\rm FR}]\simeq N\ln c$ and for an NBW on a configuration model network with a power-law degree distribution ${\mathbb{E}}[T_{\rm FR}]\sim Nc/k_{\rm min}$. In light of these results, it is interesting to note that the dependence of the mean first return time on the mean degree $c$ is a non-trivial issue, which depends on the details of the degree distribution. In the examples studied here we observe three different behaviors: in RRGs the mean first return time is independent of $c$, in ER networks it decreases with $c$ and in configuration model networks with an exponential degree distribution it increases with $c$.
@@
In all the network ensembles considered above, the long-time tail of $P(T_{\rm FR}>t)$ exhibits a decaying exponential form, which is determined by the lowest-degree nodes in the network. From a broader perspective, it implies that the distribution of first return times is mostly characterised by low-degree nodes that reside in the periphery of the network. This is unlike the outburst dynamics of other processes such as the spreading of information and infections, which are dominated by the highest degree nodes (or hubs) that reside in the core of the network [5].
@@
Apart from the first return process, there are other significant events that take place over the lifetime of an NBW (and other RWs) on a random network. One of them is the first hitting (FH) process, which is the first time at which an NBW steps into a previously visited node. Starting from a random initial node $i$, in the early stages of its trajectory, an NBW visits a new node at each time step. During this time, the statistical properties of the NBW trajectory are identical to those of a self avoiding walk [71]. After the first hitting event, in some of the time steps the NBW visits yet-unvisited nodes and in other time steps it revisits nodes that it has already visited before. The distribution of first hitting times of RWs and NBWs on ER networks were studied in Refs. [72] and [73], respectively. It was found that in both cases, for sufficiently dense ER networks in which there are no leaf nodes of degree $k=1$, the distribution $P(T_{\rm FH}>t)$ of first hitting times is given by a product of an exponential distribution and a Rayleigh distribution, which is a special case of the Weibull distribution. In this limit, the mean first hitting time of NBWs on ER networks is given by
@@
The results presented in this paper were derived in the context of configuration model networks. However, we expect them to apply within a good approximation to a somewhat broader class of small-world networks which are sufficiently strongly connected without bottlenecks and exhibit short mixing times of the NBW, determined by the spectral gap of the non-backtracking (Hashimoto) matrix. In contrast, these results are not expected to apply in the case of modular networks, which consist of several modules with weak connections between them and for networks that exhibit long mixing times. However, in the case of modular networks, at short times, the distribution of first return times is likely to behave as if the module on which the initial node resides is isolated from other modules. This behavior persists until the probability that the NBW will hop into some other module becomes significant.
@@
Another key factor influencing the mixing times is the clustering coefficient, primarily mediated through its effect on the spectral gap. The clustering coefficient measures how often a node’s neighbors form triangles indicating the degree of local inter-connectedness. Networks with higher clustering coefficients typically have smaller spectral gaps. This occurs because increased clustering introduces more local structure. A smaller spectral gap leads to longer mixing times, since random walks become trapped in tightly connected neighborhood before fully exploring the network [80]. Consequently, higher clustering increases the mixing times. We thus expect our results to be valid as long as the clustering is not too strong.